January 7, 2011

OBSERVERS: Jill Bechtold, Dave Thompson

VERY CLEAR all afternoon. Preparing Lucifer.

Changed Save path to /data/luci/20110108/

3:48pm Took darks from last night. Turned off lights in dome. Forgot to set Archive Info.

4:10pm Darks done. Joar and company are installing stuff on the telescope.

4:49pm. Opening Dome. Very clear out. Left door latch won't retract, can't open, John Little is investigating.

5:56pm. Still can't open.

7:45pm. Still working on opening, Bruce Bingham on his way from Safford. I'm learning how to make authentic empanadas.

9pm. Very clear out.

9:32pm Opening.

9:42pm. Slewing to pointing star near MS0451.

9:44pm. Going to standard to collimate.

9:54pm Executing ms0451.Kacq.txt with "Wait for collimation" ON

9:58pm. Executing ms0451.K.txt Seeing .8" 20 minute exposure. Nice! Could easily align slit, if only we could.

Wait for collimation is OFF

10:30 pm Executing ms0451.Hacq.txt

10:32pm Executing ms0451.H.txt 20 min exposure seeing .87"

10:49pm Some thin clouds -- guide star fading. Nothing obvious cloud-wise on weather map. Guide star is 15th mag

10:55pm Dither stopped, lost guide star in clouds. RATS!

Clouds. Having trouble seeing 7magn star.

11:19pm Trying W3 field for Serena. 12mag guide star. Executing W3KimKs_1.txt . See GS but it's faint. Too cloudy, can't collimate

11:30pm. Two stars outside, that's it. bummer.

12:15am. Off to Abell 697 to see if we can see guide stars there. Lots of clouds. Pointed and collimated on a VERY bright star, lost it several times.

12:25am. Executing imaging_a0697_K.txt. Seeing 1.2 arcsec. Lots of clouds. Trying 30 min exposure. Getting something.

12:53am. Script stopped because it thought we lost guide star, even though it looks OK.

12:55am. Restarted, executing restof_a0697_K.txt (no telescope preset, 30 minutes) Lost guide star after one exposure. Kept going, but not getting anything useful.

Lost guide star and script stopped after 10 exposures.

1:14am. Restarted restof_a0697_K.txt

1:25am. Restarted restof_a0697_K.txt after it bombed for losing guide star. Humidity 68% and rising.

1:36am. Lost guide star, script stopped.

1:45am. Restarted script, but then lost guide star. Humidity 76%

2:23am. Restarted for a little while, then lost guide star.

3:10am. Off to try to set up on MACS 1149. Still very cloudy. Humidity 82%.

3:20am. Starting macs1129.Kacq.txt. Sky is clearer.

3:25am. Started macs1129.K.txt. Before one exposure ends, DEWPOINT alarm went off. We have to close! Are you kidding?

3:32am. Started Darks for tonight. Humidity 88% and rising. Dome not quite closed when began darks.

3:42am. Humidity 90%

Topic revision: r3 - 08 Jan 2011, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback